# VC Search: Bridging the Gap Between Well-Defined and Ill-Defined Problems in Mathematical Reasoning

### v2

@@ 5 Related work
Mathematical Reasoning for LLMs. Mathematical reasoning is a crucial aspect in evaluating model reasoning skills, and there are currently two predominant lines for enhancing these skills. One line involves leveraging the existing few-shot prompt tool, such as CoT Wei et al. (2022), PAL Gao et al. (2023). The other is centered around fine-tuning strategy, like Metamath Yu et al. (2023), WizardMath Luo et al. (2023) and Mugglemath Li et al. (2023). Recent work has focused on how to achieve results that match or even exceed those of large models on smaller models Guan et al. (2025) and smaller training datasets Li et al. (2024a) by introducing techniques such as reinforcement learning and MCTS Tolpin and Shimony (2012).
LLM robustness. Previous work can be broadly categorized into two types, perturbations to model inputs and prompting with noisy ground truth. The generation of anticonceptual examples Jia and Liang (2017); Morris et al. (2020); Wang et al. (2021) and irrelevant context Sinha et al. (2019); Clark et al. (2020); Han et al. (2022) are the two main types of input perturbations. In the specific domain of Math Word Problems (MWP),  Zhou et al. (2024) introduces irrelevant punctuation marks as distractors, while  Shi et al. (2023) employs a sentence of unrelated contextual text to serve as a distractor, both aiming to investigate model performance variations. The work most similar to ours is MathTrap Zhao et al. (2024), which focuses on a relatively small set of fewer than 300 ill-defined problems. In contrast, our PMC dataset is far more comprehensive, containing over 5,000 ill-defined problems.
Neuro-Symbolic Methods about LLM reasoning. The primary challenge of these methods lies in ensuring that the LLM correctly translates the reasoning problem from natural language (NL) to the formal language understood by the solver Raza and Milic-Frayling (2025). For instance, Logic-LM Pan et al. (2023) utilizes LLMs to convert natural language into symbolic formulas, SatLM Ye et al. (2024) enables LLMs to generate task specifications that assist in translating natural language into logical predicates. LOT Liu et al. (2024b), similar to CoT, generates progressive logical paths. However, many of these methods struggle to extend successfully to smaller models, due to their limited contextual learning capabilities and lack of formal reasoning knowledge.
@@ 6 Conclusion
Time Consumption. Due to the use of variable-wise refinement and search architecture during the reasoning process, our method incurs higher time overhead compared to the baseline methods.
Limitations of Formal Tools. Our ability to identify ill-defined problems relies on formal tools, such as SMT solver. According to the algorithm design, the system will directly reject tasks that are unsuitable for modeling with logical tools, which may lead to the incorrect rejection of some well-defined problems.
@@ Appendix A Appendix
A.1 Details of PMC
We give more details of our PMC here. We show the number of specific subsets of PMC in Table 5, and show more representative problems to help understand our dataset.
@@ A.1 Details of PMC
A.2 Details of VCSearch
In this part we will introduce the details in our algorithm.
@@ A.2.1 Prompts in VCSearch
We show the prompts we use in VCSearch with examples 3 and 4.
@@
1. Based on the problem description, consider whether the current constraints accurately capture the problem. 2. Add constraints based on real-world knowledge, considering whether there are any missing modeling statements, such as the quantity of items should be >= 0, or the relationships between the sides of a triangle. Please note that you only need to add constraints to the CURRENT HEAD VARIABLE; in other words, the new constraints MUST include the head variable! You can first provide your thought process, and then write the new constraints that include the head variable after the identifier <SOS> You can follow the example: Question: Josh decides to try flipping a house. He buys a house for \$80,000 and then puts in \$50,000 in repairs. This increased the value of the house by 150%, but the market value of the house after repairs is only \$100,000. How much profit did he make? Existing Constraints: solver.add(initial-cost == 80000) solver.add(total-investment == initial-cost + repair-cost) solver.add(repair-cost == 50000) solver.add(increased-value-percentage == 0.5) # 150% increase solver.add(expected-value == initial-cost * (100 + increased-value-percentage)) solver.add(market-value-after-repairs == 100000) solver.add(total-investment >= 0) solver.add(profit == market-value-after-repairs - total-investment) solver.add(profit >= 0) solver.add(expected-value >= 0) Now head variable: expected-value Now existing constraints with head variable: solver.add(expected-value == initial-cost * (100 + increased-value-percentage)) solver.add(expected-value >= 0) Answer: 1. For the constraint expected-value == initial-cost * (100 + increased-value-percentage), the equation for expected-value in the problem should be initial-cost * (1 + increased-value-percentage). Therefore, this constraint should be modified to solver.add(expected-value == initial-cost * (1 + increased-value-percentage)). 2. For the constraint solver.add(expected-value >= 0) aligns with real-world requirements. Additionally, since expected-value is an unknown variable, it is appropriate to add real-world constraints, so this should be retained. 3. Furthermore, expected-value and market-value-after-repairs refer to the same entity in the problem, so a constraint should be added: market-value-after-repairs == expected-value. <SOS> So, new Constraints with head variable is solver.add(expected-value == initial-cost * (1 + increased-value-percentage)) solver.add(expected-value >= 0) solver.add(expected-value == market-value-after-repairs) Question:{question} Existing Constraints:{constraint} Now head variable:{head} Now existing constraints with head variable:{constrain-head} Answer:
@@ A.3.1 Setup
Compared methods. We selected three representative few-shot prompting methods, along with the zero-shot method that utilizes the intrinsic capabilities of the model, and compared them with our proposed VCSearch method. The methods are introduced as follows: (1)Basic, which is the zero-shot baseline method, directly feeds the problem and instructions to the LLMs without any example problem in the context. (2)CoT,  Wei et al. (2022), requires the model to explicitly output intermediate step-by-step reasoning through natural language before providing the final answer. (3)PAL Gao et al. (2023), converts each step of problem-solving into a programming language format and subsequently utilizes an external programming language interpreter for execution, thereby obtaining the results. (4)SMT, utilizes SMT-LIB to model the problems, then uses an external SMT solver to check for a feasible solution to the problem as well as obtain the ground-truth answer.
@@ A.3.2 Prompts used in Preliminary experiments
A.3.3 More experiment results
@@
Here, we also tested our method on several other benchmarks that involve refusal to answer. Our method also demonstrated superior performance on MathTrap. However, MathTrap’s mathematical problems involve a significant amount of geometry and algebra, which are not well-suited for formal tool modeling. This is also not suitable for methods such as PAL. So we only compare ours with zero-shot method. In such scenarios, our method adopts a relatively conservative approach, rejecting any problem it cannot confidently solve in order to maintain the safety of the reasoning system.

### v3

@@ 5 Related work
Enhancing Mathematical Reasoning in LLMs Mathematical reasoning is a crucial aspect in evaluating model reasoning skills Xiong et al. (2025), and there are currently two predominant lines for enhancing these skills. One line involves leveraging the existing few-shot prompt tool, such as CoT Wei et al. (2022), PAL Gao et al. (2023). The other is centered around fine-tuning strategy, like Metamath Yu et al. (2023), WizardMath Luo et al. (2023) and Mugglemath Li et al. (2023). Recent work has focused on how to achieve results that match or even exceed those of large models on smaller models Guan et al. (2025) and smaller training datasets Li et al. (2024a) by introducing techniques such as reinforcement learning and MCTS Tolpin and Shimony (2012).
Robust Mathematical Reasoning Model robustness is essential for secure deployment Sima et al. (2025), particularly in critical downstream applications such as finance Cao (2022) or healthcare Tian et al. (2025b). Traditional approaches, however, have mainly emphasized performance under noisy Liu et al. (2022), partially supervised Tian et al. (2024), or open-world settings Zhou et al. (2024a, 2025a). In recent years, there has been a significant surge in attention to the robustness of LLMs Morris et al. (2020); Wang et al. (2021). In the context of robust mathematical reasoning, most existing work focuses on defining and constructing challenging "trap" datasets. For instance, Wang et.al  Wang et al. (2024) treats mathematical problems from different datasets as an out-of-distribution (OOD) generalization problem. Robustmath Zhou et al. (2024b) introduces irrelevant punctuation marks as distractors, while GSMIC Shi et al. (2023) employs a sentence of unrelated contextual text to serve as a distractor, both aiming to investigate model performance variations. Recently, GSM-DC Yang et al. (2025a) analyzes how LLM reasoning is distracted by irrelevant context through controllable data generation. The work most similar to ours is MathTrap Zhao et al. (2024), which focuses on a relatively small set of fewer than 300 ill-defined problems. In contrast, our PMC dataset is far more comprehensive, containing over 5,000 ill-defined problems.
Neuro-Symbolic Methods with LLM reasoning. Neuro-symbolic Colelough and Regli (2025); Yang et al. (2025b) methods have recently emerged as an effective approach to enhancing model reasoning capabilities and have been widely applied to reasoning augmentation and data generation across various downstream domains, including math Mirzadeh et al. (2024), law Zhou et al. (2025b) and tabular data Tian et al. (2025a). The primary challenge of these methods lies in ensuring that the LLM correctly translates the reasoning problem from natural language (NL) to the formal language understood by the solver Raza and Milic-Frayling (2025). For instance, Logic-LM Pan et al. (2023) utilizes LLMs to convert natural language into symbolic formulas. SatLM Ye et al. (2024) enables LLMs to generate task specifications that assist in translating natural language into logical predicates. LOT Liu et al. (2024b), similar to CoT, generates progressive logical paths. However, many of these methods struggle to extend successfully to smaller models, due to their limited contextual learning capabilities and lack of formal reasoning knowledge.
@@ 6 Conclusion
Time Consumption. Due to the use of variable-wise refinement and a search-based architecture during the reasoning process, our method inevitably incurs higher computational overhead compared to baseline approaches. While this additional cost is the price for improved robustness and broader applicability, it may limit scalability when applied to very large datasets or real-time scenarios. Future work may investigate techniques for reducing this overhead, such as pruning strategies or parallelization.
Limitations of Formal Tools. Our method’s ability to identify ill-defined problems heavily relies on formal tools, such as SMT solvers. By design, the system will directly reject tasks that cannot be adequately modeled with logical constraints. Although this ensures rigor in handling pathological cases, it may also lead to overly conservative behavior, including the incorrect rejection of certain well-defined problems. Extending the framework with more flexible or hybrid reasoning mechanisms could help alleviate this limitation.
Acknowledgements
This research was supported by the Jiangsu Science Foundation (BG2024036, BK20243012), National Natural Science Foundation of China (624B2068,62576162,62576174), and the Fundamental Research Funds for the Central Universities (022114380023).
@@ Appendix A Appendix
The appendix is organized as follows: Section A.1 provides additional details of our proposed dataset PMC; Section A.2 describes the operation process of our algorithm VCSearch; and Section A.3 presents further experiments and discussions.
A.1 Details of benchmark PMC
We give more details of our constructed benchmark PMC here.
A.1.1 Composition and examples of PMC
We show the number of specific subsets of PMC in Table 5, and show more representative problems to help understand our dataset.
@@ A.1.1 Composition and examples of PMC
A.1.2 Constrction prompt
The construction prompt we used is shown in the example 3,4,5.
Example 3: Constrction prompt for missing type Given the following math problem, identify all the variables and constraints involved. Then, modify the problem by replacing a key numerical value in one of the constraints with an indefinite placeholder (e.g., “some number”, “a certain value”, etc.), such that the resulting problem lacks sufficient information to determine a unique solution. You can answer with following step: Step 1: Variable and Constraint Identification. Step 2: Decide the mutated Variable or constraint and explain the reason. Step 3: Answer with final mutated problem. Original Problem: {Problem} Modified Problem: [Your answer]
Example 4: Constrction prompt for contra type Given the following math problem, identify all the variables and constraints involved. Then, modify the problem by introducing an additional constraint that directly conflicts with an existing one. The resulting problem should contain contradictory information that makes it logically unsolvable. You can answer with following step: Step 1: Variable and Constraint Identification. Step 2: Decide the mutated Variable or constraint and explain the reason. Step 3: Answer with final mutated problem. Original Problem: {Problem} Modified Problem: [Your answer]
Example 5: Validation prompt Given the following math problem, determine whether it is solvable. If not, identify why the problem is ill-defined. Specifically, analyze whether the conditions provided are insufficient or self-contradictory, making it impossible to derive a unique solution. You can answer with the following steps: Step 1: Variable and Constraint Identification. Step 2: Analyze whether the problem is solvable under the given constraints. If it is unsolvable, explain whether it is due to missing information or contradictory conditions, and identify the responsible part(s). Step 3: Give the final feedback if the question is unsolvable Problem: {Problem} Answer: [Your answer]
A.1.3 Human annotators
When the LLM used for verification outputs inconsistent responses, we will enable human annotators to verify. Our annotators come from within the lab, no more than 5 master’s and doctoral students.
A.2 Details of algorithm VCSearch
In this part, we will introduce the details of our algorithm VCSearch.
@@ A.2.1 Prompts in VCSearch
We show the prompts we use in VCSearch with examples 6 and 7.
@@
1. Based on the problem description, consider whether the current constraints accurately capture the problem. 2. Add constraints based on real-world knowledge, considering whether there are any missing modeling statements, such as the quantity of items should be >= 0, or the relationships between the sides of a triangle. Please note that you only need to add constraints to the CURRENT HEAD VARIABLE; in other words, the new constraints MUST include the head variable! You can first provide your thought process, and then write the new constraints that include the head variable after the identifier <SOS> You can follow the example: Question: Josh decides to try flipping a house. He buys a house for \$80,000 and then puts in \$50,000 in repairs. This increased the value of the house by 150%, but the market value of the house after repairs is only \$100,000. How much profit did he make? Existing Constraints: solver.add(initial-cost == 80000) solver.add(total-investment == initial-cost + repair-cost) solver.add(repair-cost == 50000) solver.add(increased-value-percentage == 0.5) # 150% increase solver.add(expected-value == initial-cost * (100 + increased-value-percentage)) solver.add(market-value-after-repairs == 100000) solver.add(total-investment >= 0) solver.add(profit == market-value-after-repairs - total-investment) solver.add(profit >= 0) solver.add(expected-value >= 0) Now head variable: expected-value Now existing constraints with head variable: solver.add(expected-value == initial-cost * (100 + increased-value-percentage)) solver.add(expected-value >= 0) Answer: 1. For the constraint expected-value == initial-cost * (100 + increased-value-percentage), the equation for expected-value in the problem should be initial-cost * (1 + increased-value-percentage). Therefore, this constraint should be modified to solver.add(expected-value == initial-cost * (1 + increased-value-percentage)). 2. For the constraint solver.add(expected-value >= 0) aligns with real-world requirements. Additionally, since expected-value is an unknown variable, it is appropriate to add real-world constraints, so this should be retained. 3. Furthermore, expected-value and market-value-after-repairs refer to the same entity in the problem, so a constraint should be added: market-value-after-repairs == expected-value. <SOS> So, new Constraints with head variable is solver.add(expected-value == initial-cost * (1 + increased-value-percentage)) solver.add(expected-value >= 0) solver.add(expected-value == market-value-after-repairs) Question:{question}
Existing Constraints:{constraint}
Now head variable:{head}
Now existing constraints with head variable:{constrain-head}
Answer:
@@ A.2.3 Double-check solving strategy with SMT solver
A.2.4 An example for VCSearch
Our approach to determining variable-constraint relationships is as follows:
  • Preparation Phase (Variables → Constraints): For a given variable, directly retrieve all constraints containing that variable from the constraint pool.
  • Update Phase (Constraints → Variables): For a given constraint, we identify all new associated variables in it.
To further illustrate this method, we present a concrete example using a contra-type problem in PMC (example 8) to demonstrate the search process:
Example 8: Example in VCSearch "Josh decides to try flipping a house. He buys a house for 80,000 and then puts in 50,000 in repairs. This increased the value of the house by 150%, but the market value of the house is only \$100,000. How much profit did he make?" After the initialization step, we obtain an initial constraint system, represented in Python Z3 code. This system consists of a variable queue and a constraint pool. Variables: "initial-cost", "repair-cost", "increased-value-percentage", "expected-value", "market-value-after-repairs", "profit", "total-investment" Constraints:
initial-cost == 80000
repair-cost == 50000
market-value-after-repairs == 100000
increased-value-percentage == 0.5
total-investment == initial-cost + repair-cost
expected-value == initial-cost * (100 + increased-value-percentage)
profit == market-value-after-repairs - total-investment
After the Initialization, assume that the first element in the variable queue is "expected-value", we will demonstrate a single iteration of the search process. Preparation Identify constraints involving this variable "expected-value":
expected-value == initial-cost * (100 + increased-value-percentage)
Exploration Utilize LLM knowledge to refine the constraints by generating a constraints set with the head variable "expected-value":
expected-value == initial-cost * (basic_multiplier + increased-value-percentage)
basic_multiplier == 1
Verification Compare the original constraint system with the refined one and select the better version. (In this case, the newly generated constraint set is selected). Update Replace the outdated constraint with the refined one. Identify any newly introduced variables (e.g., "basic_multiplier") and append them to the tail of the variable queue for subsequent iterations.
@@ A.3 Details of Experiment
In this section, we provide additional experimental details and discussions. However, due to time and computational constraints, some of the analyses were conducted on a subset of the dataset.
@@ A.3.1 Setup
Compared methods. We selected three representative few-shot prompting methods, along with the zero-shot method that utilizes the intrinsic capabilities of the model, and compared them with our proposed VCSearch method. The methods are introduced as follows: (1)Basic, which is the zero-shot baseline method, directly feeds the problem and instructions to the LLMs without any example problem in the context. (2)CoT,  Wei et al. (2022), requires the model to explicitly output intermediate step-by-step reasoning through natural language before providing the final answer. (3)PAL Gao et al. (2023), converts each step of problem-solving into a programming language format and subsequently utilizes an external programming language interpreter for execution, thereby obtaining the results. (4)Satlm Ye et al. (2024), utilizes SMT-LIB to model the problems, then uses an external SMT solver to check for a feasible solution to the problem as well as obtain the ground-truth answer.
@@ A.3.2 Prompts used in Preliminary experiments
A.3.3 More experiment results on other benchmark
@@
Here, we also tested our method on several other benchmarks that involve refusal to answer. Our method also demonstrated superior performance on MathTrap. However, MathTrap’s mathematical problems involve a significant amount of geometry and algebra, which are not well-suited for formal tool modeling. This is also not suitable for methods such as PAL. So we only compare ours with the zero-shot method. In such scenarios, our method adopts a relatively conservative approach, rejecting any problem it cannot confidently solve in order to maintain the safety of the reasoning system.
A.3.4 More ablation about VCSearch
We further conduct ablation of the algorithm from the following aspects:
Variable Ordering. By default, we refine variables in the order they appear in the problem statement. This approach often approximates the topological order of problem-solving steps, especially when the number of solution steps is limited. We also experimented with an alternative order: refining variables based on their frequency of occurrence in the constraints (from highest to lowest).
As shown in the table below, using the frequency-based iteration order improved performance on "contra" type problems while degrading it on "missing" type problems, with "well-defined" problems remaining stable. Our analysis suggests this is because "contra" problems often have contradictions embedded in hidden variables at constraint intersections (with high occurrences), which are more readily identified and optimized with frequency-based ordering. Conversely, in "missing" type problems, the missing variables frequently appear earlier in the data, making sequential iteration more effective.
Iteration Strategy. We experimented with the number of variable iterations as a hyperparameter T. While additional iterations bring slight performance improvements, they also introduce significant increases in computational and time costs. After weighing performance gains against resource consumption, we adopted a single-iteration setting as the most cost-effective choice. Detailed results are presented in the table below. (Due to time and computational limitations, we report results using the Qwen-7B model, based on 300 randomly sampled instances per dataset.)
Table 7: Performance under different configurations
Default (T=1, appear-order) T=2 T=3 frequency-order
Contra 56.89 57.67 56.33 61.00
Missing 92.60 93.00 91.67 87.67
Well-type 79.09 77.00 79.67 79.00
A.3.5 Discussion about reasoning in realistic scenarios
Discussion of dataset ratios
In our paper, we adopted a balanced setting($i.e.,D_{w}:D_{i}=1:1$) to measure the reaction score. This balanced approach allows us to evaluate the capability of methods to both answer well-defined problems and reject ill-defined problems with equal importance. This evaluation strategy is analogous to how imbalanced classification studies often report balanced metrics to properly assess model performance across all classes Thabtah et al. (2020). By maintaining this balanced setting, we provide a more comprehensive and fair assessment of each method’s capabilities of answering and rejecting. Additionally, we compared the R-score performance across different dataset ratios (defined as $\alpha=D_{w}:D_{i}$) on the Qwen1.5B model, and our method consistently demonstrated superior results.
Table 8: Performance among different data ratios
$\alpha$ 0.2 0.5 1 2 5
CoT $44.61\pm 1.02$ $49.58\pm 2.00$ $53.91\pm 1.16$ $58.96\pm 0.78$ $62.83\pm 1.55$
CoT + Ours $64.40\pm 0.43$ $64.05\pm 0.60$ $63.26\pm 0.84$ $64.33\pm 0.89$ $62.91\pm 0.79$
PAL $16.01\pm 0.66$ $24.03\pm 1.12$ $32.85\pm 1.00$ $41.15\pm 0.49$ $49.53\pm 2.89$
PAL + Ours $65.26\pm 1.54$ $62.46\pm 0.22$ $62.28\pm 0.65$ $58.55\pm 1.13$ $58.84\pm 0.56$
More convincing metrics
To prevent excessive score inflation through question rejection (where rejecting all questions would yield only 50% of the total score), we introduce the R*-score metric as below
$\frac{\sum_{p\in D_{i}}\mathbb{I}[f(p)=\text{Reject}]+\sum_{p\in D_{w}}\mathbb{I}[f(p)=g(p)]}{|D_{i}|+|D_{w}|}$
We evaluate our method under balanced settings and present the results in the following table. Our approach maintains superior performance in most scenarios(R*-score), demonstrating that our performance gains do not stem from simply rejecting most questions.
Table 9: Perfomance among R-score and R*-score
Qwen 1.5B Qwen 3B
Method R-score R*-score R-score R*-score
CoT $53.91\pm 1.16$ $51.10\pm 2.08$ $65.93\pm 0.73$ $65.10\pm 1.04$
CoT + Ours $63.26\pm 0.84$ $53.10\pm 0.06$ $73.98\pm 0.28$ $66.93\pm 0.28$
PAL $32.85\pm 1.00$ $30.63\pm 0.18$ $48.56\pm 0.22$ $47.66\pm 0.49$
PAL + Ours $62.28\pm 0.65$ $51.90\pm 1.15$ $74.08\pm 1.17$ $65.73\pm 1.30$
A.3.6 Performance on Larger LLM
We’ve extended our experimental results to include GPT-4-0613 and GLM-4-Plus. These findings indeed demonstrate that larger language models exhibit a stronger ability to identify ill-defined problems. However, it’s crucial to highlight that even with these powerful models, our proposed method can further enhance their capability to recognize and handle such issues, significantly improving their robustness, particularly in "contra" type problems.
Model Dataset Basic Ours
gpt4-0613 Contra 35.00 71.00
Missing 82.00 85.00
GLM-4-Plus Contra 44.67 64.00
Missing 73.67 86.00
Table 10: Performance comparison between baseline and our method.
Even though the ability of large LLMs to handle robust mathematical reasoning has improved, this doesn’t diminish the importance of addressing ill-defined problems. In reality, not all application scenarios possess the resources or infrastructure to deploy ultra-large models. For instance, mid-sized models (around 7 billion parameters) are widely adopted in practical applications due to their excellent deployability and cost-effectiveness. In these contexts, robust reasoning capabilities remain a critical focus.
Furthermore, we contend that symbolic methods offer a unique advantage in this scenario, rather than being overly complex. Specifically, a symbolic solver can be effectively utilized as a tool to assess the completeness of problem descriptions, while the language model focuses on modeling the constraint system. This division of labor frees the model from the trade-off between identifying pathological problems and solving well-defined ones. And even with advanced larger LLMs, experiments show they can’t fully escape this trade-off, yet our symbolic approach still proves effective.
A.3.7 Computational cost discussion
Our framework employs a sequential iterative structure. In terms of memory and computational costs, it’s comparable to standard LLM inference, as most resource consumption doesn’t significantly increase. The primary overhead lies in runtime. We use the Z3 SMT solver, which has a relatively low CPU footprint. For example, when running Qwen-7B on our proposed PMC dataset with an RTX 4090, GPU and CPU memory consumption are approximately 15 GB and 1.6 GB, respectively.
While our method does have a higher runtime than zero-shot/few-shot inference, this is a common characteristic of test-time scaling approaches. Our proposed method achieves linear time complexity with respect to the number of variables, rather than exponential growth. Our empirical runtime measurements confirm this linear scaling behavior, significantly outperforming methods that build constraint systems using tree search. The table below shows the average time required for the Qwen-7B model to solve a single problem using different methods.
Method Basic SMT Ours Tree Search
Time 6.6s 12s 79.8s 156s
Table 11: Time consumption comparison of different methods.
A.3.8 LLM verification in VCSearch
We acknowledge the current limitations of Large Language Models (LLMs) in modeling symbolic systems. However, it’s important to clarify that our LLM-Judge relies not solely on the model’s output. To enhance the reliability of its judgments, we also provide the problem’s text description and the SMT solver’s execution results (Equation 7) as additional inputs.
Given that incorporating human evaluation into every LLM-Judge process would demand substantial annotation resources and time, it’s impractical for real-world application. Therefore, we propose an alternative metric: Judge-Error-Rate (JER). Among samples where the final output was incorrect, any correct answer that was found during the search process but not ultimately selected as the final output is counted as a judge error. We calculate JER as the proportion of these judge errors among all judging instances. This metric serves as an effective measure of LLM-Judge’s reliability. We’ve calculated the JER for our method across various models and datasets, and the results demonstrate the high effectiveness of our LLM-Judge approach.
Table 12: JER on different datasets
Model Contra Missing Well-defined
Qwen7b 8.80 1.11 1.69
Qwen3b 11.64 2.60 1.68
Qwen1.5b 1.40 9.26 2.44